# All-in-one: Understanding and Generation in Multimodal Reasoning with the MAIA Benchmark

### v2

@@ 4.3 Baselines
This baseline applies only to Task 1 and selects the most probable statement in a TS-FS pair. It serves as a unimodal language baseline, reflecting the distributional biases of LLMs. Probabilities of TS and FS are first estimated on five open-weight LLMs that have shown good performance on a variety of Italian tasks Magnini et al. (2025): Llama-3.1 ($8$B-Instruct), LLaMAntino-2 ($7$B), LLaMAntino-3-ANITA ($8$B-Instruct), Gemma ($7$B) and Qwen2.5 ($7$B-Instruct). For each TS-FS pair, we selected the item with the highest probability among the five models.
@@
replaces MAIA videos with a fully black clip, used as a proxy for a no-video condition This setup minimizes access to visual features, pushing the model to rely mainly on the language component, while the true unimodal evaluation remains the previous one.
@@
4.4 Vision-Language-Models.
We benchmarked four recent VLMs sourced from the Hugging Face Hub, representing state-of-the-art approaches in Vision-Language tasks: [^9] More details about VLMs are reported in Appendix B.1 InternVL2 (8B, Chen et al. (2024)), LLaVA-NeXT-Video (7B, Zhang et al. (2024b)), LLaVa-OneVision (7B, Li et al. (2024a)), and Qwen2.5-VL (7B, Qwen et al. (2025)). All models accept a [video, text] pair as input, and uniformly sample 32 frames from the video.
@@ 5.1 Results on Visual Statement Verification
Table 1 shows VLM accuracy across the three settings (black video, 1-Frame, 32-Frames) and reasoning categories when models consistently answers correctly (i.e., by choosing 8/8 the TSs within the pool). Qwen2.5-VL achieves the highest accuracy in the 32-Frame setup with an average score of 0.44, marking an 8-point improvement over the correspondent 1-Frame settting. Llava-Next-Video shows the weakest performance across all three configurations, likely due to its underlying Vicuna-7B LLM, weakly trained in Italian. All models outperform the unimodal baseline (0.05): in the 32-Frame setting, gains go from +31 points (InternVL2) to +44 (Qwen2.5-VL), confirming the use of visual cues to counteract language-driven biases.
Figure 3 helps visualizing the comparison across reasoning categories. Here we report results only for Qwen-2.5-VL, the best model, while similar figures for the other models are in Appendix B (Figure 15, 16 and 19). The most difficult categories are Counterfactual, Planning, Implicit Partial, Spatial Partial, and Temporal-Duration. On the other side, Causal, Sentiment, Implicit Partial and Total, and Spatial Partial and Total are the categories for which the model profits the most from visual clues, by leveraging the broader visual window provided in the 32-Frame setting.
@@
Figure 3a vs.  3b highlights the role of the consistency check in MAIA by using a pool of 8 TS-FS pairs. When pairs are considered independently, as it is usually done in VLM benchmarks, the model’s performance increases significantly, showing that it relies on spurious correlation, effects that is strongly mitigated by the MAIA’s severe evaluation regime. Such trend is even much more visible in the black video setting than in the 32-Frames one (e.g. +51 vs. +40 for Qwen2.5-VL). By systematically comparing models’ performance in the independent and pool-aggregated settings, we see that this is a general finding across models. Moreover, we find that Qwen-2.5-VL is not only the best-performing model, but also the most consistent, with a lower drop in accuracy, particularly in the 32-Frames setting (Details in Appendix B.2).
@@ 5.2 Results on Open-ended Generation
Black video InternVL2 0.37 0.42 0.60 0.30 0.68 0.43 0.08 0.21 0.23 0.36 0.25 0.55 0.25
Llava-Next-Video 0.27 0.43 0.47 0.30 0.40 0.33 0.51 0.21 0.12 0.16 0.04 0.12 0.16
Llava-oneVision 0.40 0.66 0.68 0.29 0.60 0.60 0.28 0.26 0.24 0.36 0.18 0.37 0.23
Qwen-2.5-VL 0.35 0.36 0.69 0.08 0.54 0.74 0.23 0.24 0.19 0.30 0.19 0.41 0.20
1-Frame InternVL2 0.44 0.57 0.65 0.21 0.65 0.60 0.10 0.33 0.38 0.39 0.47 0.53 0.35
Llava-Next-Video 0.32 0.30 0.56 0.20 0.46 0.60 0.29 0.18 0.32 0.24 0.27 0.20 0.19
Llava-oneVision 0.50 0.59 0.78 0.15 0.66 0.74 0.37 0.39 0.54 0.39 0.51 0.54 0.33
Qwen-2.5-VL 0.51 0.56 0.76 0.32 0.60 0.79 0.40 0.35 0.50 0.38 0.53 0.55 0.38
32-Frames InternVL2 0.49 0.54 0.68 0.28 0.64 0.62 0.11 0.45 0.48 0.47 0.51 0.57 0.46
Llava-Next-Video 0.33 0.37 0.38 0.16 0.42 0.48 0.27 0.24 0.37 0.29 0.39 0.32 0.29
Llava-oneVision 0.53 0.67 0.79 0.11 0.65 0.79 0.23 0.55 0.56 0.51 0.62 0.40 0.46
Qwen-2.5-VL 0.61 0.71 0.80 0.43 0.60 0.85 0.55 0.55 0.60 0.50 0.70 0.54 0.53
Table 2: OEVQA (Task $2$): accuracy of correct answers with LLM-as-a-judge.
@@
Black video Qwen-2.5-VL 0.09 0.05 0.29 0.05 0.09 0.14 0.02 0.02 0.20 0.01 0.03 0.10 0.04
1-Frame Qwen-2.5-VL 0.22 0.25 0.23 0.25 0.11 0.40 0.33 0.16 0.28 0.06 0.28 0.17 0.14
32-Frames Qwen-2.5-VL 0.31 0.41 0.26 0.31 0.16 0.48 0.41 0.31 0.38 0.19 0.44 0.22 0.19
Table 3: Aggregate accuracy on Task $1$ (NLU) and Task $2$ (NLG) (consistency and robustness) on Qwen-2.5-VL across reasoning categories.
Table 2 reports the results on Task 2. The best performing model is again Qwen2.5-VL, reaching 0.61 accuracy in the 32-Frames setting. Hardest categories are Sentiment, Implicit Partial, Spatial Partial, Temporal, and Planning. Figure 3c shows that model profits from the 32 frames for all categories, and poorly performs when provided only with black videos, in almost all cases. Surprisingly, for Sentiment and Counterfactual, a relatively high accuracy is obtained already with black videos. Interestingly, differences also emerge when comparing the 1-Frame vs. 32-Frames settings. For Temporal Partial, and Implicit Partial, and for Spatial Total, the 1-Frame is not enough, while 32 frames increase performance. This difference is less pronounced or does not hold for Temporal Duration, Counterfactual, Planning, and Sentiment. For example, in the Implicit Partial category, the 32-Frame model answers correctly to How the man in the flannel shirt wears his hair? with The man in the flannel shirt has shaved head, while the 1-Frame setting hallucinates with The man in the flannel shirt keeps his hair in a braid.
@@
Table 4: VLM performance for OEVQA (Task 2) according to similarity-based metrics.
Table 4 highlights how similarity-based metrics (e.g., ROUGE, BLEU) often do not align with semantic correctness. InternVL2 scores highest on surface-level similarity (e.g., ROUGE:$0.61$, BERTScore:$0.84$) but lower when considering the LLM-as-a-judge metric ($0.49$), while Qwen2.5-VL offers more (semantic) accurate answers. LLaVA-NeXT-Video underperforms across all evaluations.
@@ 5.3 Discussion
The results presented for Task 1 and 2 clearly highlight two key findings. First, within each task, the role of the information extracted from videos is unequally distributed across reasoning categories. The star-shaped Figure 3b illustrates such differences: the star’s picks highlight the categories that profit from the visual input the most: Spatial Total and Partial, Total Implicit, Causal, Out-of-scope, and Sentiment. On the other hand, and quite surprisingly, Qwen2.5-VL handles Counterfactual and Uncertainty better when provided the black videos, which are expected to be uninformative, than with the full 32 frames. This calls for a deeper analysis of the reason behind such a result. Overall, this shows that MAIA categories are extremely useful to factorize the contribution of the visual vs. linguistic components of VLMs, favoring a more nuanced analysis of their actual abilities.
A second noteworthy fact is the effect of the task design. By comparing Table 1 and Table 2, we see that overall accuracy is higher in Task 2 than in Task 1, even though the underlying information the model has to exploit to perform the tasks is the same, given the alignment between their data points. Such an increase is found across models: +0.2 increase on average – in line with the Generative AI Paradox – they are better at generating than at understanding text West et al. (2024). Figure 3b vs. Figure 3b illustrates such a difference for Qwen2.5-VL. Interestingly, the difference is more pronounced for some categories, as shown by the fact that Figure 3c no longer has the shape of a star; for instance, Planning, Counterfactual, and Spatial Partial improve the most. On the other hand, Uncertainty and Out-of-scope accuracy drops in the NLG task; from a qualitative analysis, we saw that this is mostly due to the generation of hallucinations (Appendix B.4).
@@ 6 Aggregating Understanding and Generation
where $M$ is the model, $q$ a question, $S_{q}$ the set of TS-FS pairs, and $a_{M}$ the model’s answers. Intuitively, given a question $q$ on a video $v$, we reward the ability of a VLM to both select a correct answer $TF$ from a $TF-FS$ pair related to $q$, and to generate a correct answer $a$ to $q$, as discussed in Section 1 when commenting the question What is the cake made of? and the aligned data of the NLU and NLG tasks in Figure 1.
Table 3 reports Agg-Acc for our best model Qwen2.5-VL: 0.31 with 32-frames, 0.22 with 1-frame, and only 0.09 with black video; this shows the more challenging nature of the aggregate task, and that a single frame is not that robust, except for Counterfactual where reasonable answers can be produced without visual features. Planning, Spatial Partial, and Temporal Duration remain challenging even with 32-Frames. Results confirm that MAIA’s aggregated understanding&generation task creates a harder benchmark, laying the ground toward a more objective VLM evaluation.
@@ 7 Conclusion
We introduce MAIA, a benchmark designed for fine-grained investigation of the reasoning abilities of VLMs on videos. MAIA has two aligned tasks: a visual statement verification task (NLU), and a open-ended visual question answering task (NLG), both on the same set of video related questions. First, we provided a in-depth analysis of the two tasks independently, showing the importance of evaluating model with answer-pools to account for model consistency. Then, we couple comprehension and generation in an aggregated evaluation framework, arguing that the aggregated "all-in-one" understanding&generation task is a challenging and natural setting toward a more objective VLM evaluation. As for the future, it would be interesting to see whether our framework promote models that undergo learning paradigms tightly integrating these two capabilities, as in Gul and Artzi (2024).
@@
We acknowledge that the number of videos in MAIA may seem relatively small, with only 100 samples. However, their combination with 12 reasoning categories results in 19,200 samples for Task 1 and an equal number of question-answer pairs, providing a robust evaluation set, not intended for any kind of training. Still, these 100 videos constitute only the initial core of a broader evaluation framework planned for future development. Moreover, we are aware that our most probable baseline, constructed using probabilities derived from our set of LLM’s logits, poses a limitation that we intend to address in future work. We plan to compare each VLM with its corresponding LLM to obtain more reliable results for proper comparisons and analyses of potential statistical biases. Regarding Task $2$, we aim to dive deep into the comparison between the similarity metrics used and the emerging topic of LLMs as judges. Additionally, we intend to further investigate this latter approach to assess its actual validity as a reliable evaluation method. This direction will help us refine accuracy metrics, ultimately enhancing our ability to rigorously test model robustness and consistency across our two specular tasks. Furthermore, we acknowledge that our evaluation did not include large-scale propri- etary models (e.g., ChatGPT). Our focus was pri- marily on testing the performance of open-source and easily exploitable language models to provide a comprehensive overview of their capabilities on the benchmark. Nevertheless, extending our evalu- ation to bigger models would be beneficial to gain a broader understanding of their performance and to contextualize our findings within the wider land- scape of LLM research. Finally, we are also aware that our benchmark has not yet been compared with existing ones to assess its relative difficulty and challenge level. This limitation comes from the lack of comparable resources in Italian. As a future direction, we plan to translate MAIA into English and replicate our experiments, enabling more meaningful comparisons with similar English-language benchmarks.

### v3

@@ 4.3 Baselines
This baseline applies only to Task 1 and selects the most probable statement in a TS-FS pair. It serves as a unimodal language baseline, reflecting the distributional biases of LLMs. Probabilities of TS and FS are first estimated on five open-weight LLMs that have shown good performance on a variety of Italian tasks Magnini et al. (2025): Llama-3.1 (88B-Instruct), LLaMAntino-2 (77B), LLaMAntino-3-ANITA (88B-Instruct), Gemma (77B) and Qwen2.5 (77B-Instruct). For each TS-FS pair, we selected the item with the highest probability among the five models.
@@
It replaces MAIA videos with a fully black clip, used as a proxy for a no-video condition. This setup minimizes access to visual features, pushing the model to rely mainly on the language component, while the true unimodal evaluation remains the previous one.
@@
4.4 Vision-Language-Models
We benchmarked six recent VLMs sourced from the Hugging Face Hub, representing state-of-the-art approaches in Vision-Language tasks: [^11] More details about VLMs are reported in Appendix B.1 InternVL2 (8B, Chen et al. (2024)), InternVL3 (78B, Zhu et al. (2025)) LLaVA-NeXT-Video (7B, Zhang et al. (2024b)), LLaVa-OneVision (7B, Li et al. (2024a)), and Qwen2.5-VL (both 7B and 72B, Qwen et al. (2025)). All models accept a [video, text] pair as input, and uniformly sample 32 frames from the video. [^12] During the experiments with InternVL3, we found that, for about ten videos, $32$ frames exceeded the model’s context capacity, and in these cases we reduced them to $16$ to ensure proper processing.
@@ 5.1 Results on Visual Statement Verification
Table 1 shows VLM accuracy across the three settings (black video, 1-Frame, 32-Frames) and reasoning categories when models consistently answers correctly (i.e., by choosing 8/8 the TSs within the 8-item pool). Qwen2.5-VL 72B achieves the highest accuracy in the 32-Frames setup with an average score of $0.54$, marking a 14-point improvement over the correspondent 1-Frame setting. Llava-Next-Video shows the weakest performance across all three configurations, likely due to its underlying Vicuna-7B LLM, weakly trained in Italian. All other models outperform the unimodal baseline ($0.05$): in the 32-Frames setting, gains go from +$26$ points (InternVL2) to +$49$ (Qwen2.5-VL 72B), confirming the use of visual cues to counteract language-driven biases. Notably, InternVL3 78B, despite being a high-performing model, exhibits an opposite trend in the 32-Frames setting, where its accuracy drops by $0.05$ and $0.17$ compared to the Black Video and 1-Frame configurations, instead of improving as observed for the other models. [^13] One possible explanation is that the model has not been trained to process 32-frame inputs at the resolution we provide.
As a complement, Figure 3 helps visualizing the comparison across reasoning categories. Here we report results only for Qwen-2.5-VL 72B, our best model, while similar figures for the other models are in Appendix B (Figure 6, 7 and 8). The most difficult categories are Counterfactual, Planning, Implicit Partial, Spatial Partial, and Temporal Duration. In addition, Causal, Sentiment, Implicit Partial and Total, and Spatial Partial and Total are the categories for which the model profits the most from visual clues, by leveraging the broader visual window provided in the 32-Frames setting, as shown by the larger area covered by the red curve compared to the green and blue ones.
@@
Figure 3a vs.  3b highlights the role of the consistency check in MAIA by using a pool of 8 TS-FS pairs. When pairs are considered independently, as it is usually done in VLM benchmarks, the model’s performance increases significantly, showing that it relies on spurious correlation, effects that is strongly mitigated by the MAIA’s severe evaluation regime. In our case, such trend is even much more visible in the black video setting than in the 32-Frames one with a gain of +$47$ vs. +$34$ for Qwen2.5-VL 72B (see Appendix, Table 5). By systematically comparing models’ performance in the independent and pool-based settings, we see that this is a general finding across models. Moreover, we find that Qwen-2.5-VL 72B is not only the best-performing model, but also the most consistent, with a lower drop in the pool-based accuracy (Table 1) with reference to the indipendent one (Table 5), particularly in the 32-Frames setting (i.e., 34 points).
@@ 5.2 Results on Open-ended Generation
Black video InternVL2 8B 0.37 0.42 0.60 0.30 0.68 0.43 0.08 0.21 0.23 0.36 0.25 0.55 0.25
InternVL3 78B 0.52 0.77 0.96 0.00 0.83 0.74 0.09 0.33 0.35 0.50 0.54 0.49 0.62
Llava-Next-Video 7B 0.27 0.43 0.47 0.30 0.40 0.33 0.51 0.21 0.12 0.16 0.04 0.12 0.16
Llava-oneVision 7B 0.40 0.66 0.68 0.29 0.60 0.60 0.28 0.26 0.24 0.36 0.18 0.37 0.23
Qwen-2.5-VL 7B 0.35 0.36 0.69 0.08 0.54 0.74 0.23 0.24 0.19 0.30 0.19 0.41 0.20
Qwen-2.5-VL 72B 0.38 0.36 0.73 0.97 0.86 0.12 0.90 0.15 0.17 0.14 0.03 0.04 0.04
1-Frame InternVL2 8B 0.44 0.57 0.65 0.21 0.65 0.60 0.10 0.33 0.38 0.39 0.47 0.53 0.35
InternVL3 78B 0.68 0.88 0.97 0.44 0.91 0.83 0.56 0.55 0.70 0.57 0.77 0.44 0.56
Llava-Next-Video 7B 0.32 0.30 0.56 0.20 0.46 0.60 0.29 0.18 0.32 0.24 0.27 0.20 0.19
Llava-oneVision 7B 0.50 0.59 0.78 0.15 0.66 0.74 0.37 0.39 0.54 0.39 0.51 0.54 0.33
Qwen-2.5-VL 7B 0.51 0.56 0.76 0.32 0.60 0.79 0.40 0.35 0.50 0.38 0.53 0.55 0.38
Qwen-2.5-VL 72B 0.70 0.80 0.97 0.75 0.92 0.75 0.64 0.50 0.74 0.54 0.75 0.40 0.65
32-Frames InternVL2 8B 0.49 0.54 0.68 0.28 0.64 0.62 0.11 0.45 0.48 0.47 0.51 0.57 0.46
InternVL3 78B 0.77 0.86 0.96 0.54 0.77 0.85 0.64 0.73 0.81 0.50 0.54 0.49 0.62
Llava-Next-Video 7B 0.33 0.37 0.38 0.16 0.42 0.48 0.27 0.24 0.37 0.29 0.39 0.32 0.29
Llava-oneVision 7B 0.53 0.67 0.79 0.11 0.65 0.79 0.23 0.55 0.56 0.51 0.62 0.40 0.46
Qwen-2.5-VL 7B 0.61 0.71 0.80 0.43 0.60 0.85 0.55 0.55 0.60 0.50 0.70 0.54 0.53
Qwen-2.5-VL 72B 0.81 0.93 0.99 0.72 0.94 0.87 0.59 0.74 0.84 0.71 0.91 0.72 0.79
Table 2: OEVQA (Task 22): accuracy of correct answers with LLM-as-a-judge.
@@
Black video Qwen-2.5-VL 72B 0.18 0.08 0.30 0.69 0.15 0.01 0.84 0.02 0.02 0.00 0.00 0.00 0.00
1-Frame Qwen-2.5-VL 72B 0.33 0.39 0.32 0.59 0.20 0.32 0.53 0.27 0.43 0.15 0.42 0.09 0.24
32-Frames Qwen-2.5-VL 72B 0.47 0.64 0.41 0.61 0.37 0.56 0.49 0.48 0.58 0.35 0.62 0.10 0.40
Table 3: Aggregate accuracy on Task 11 (NLU) and Task 22 (NLG) (consistency and robustness) on Qwen-2.5-VL 72B across reasoning categories.
Table 2 reports the results on Task 2. The best performing model is again Qwen2.5-VL 72B, reaching $0.81$ accuracy in the 32-Frames setting. Unlike in Task $1$, here InternVL3 78B shows a positive incremental trend, with performance progressively improving from Black Video to 1-Frame and reaching its best results with 32 frames. In the latter configuration, across models, the hardest categories are Uncertainty, Out-of-scope, Implicit Partial, Temporal Duration and Partial, and Spatial Partial, while Counterfactual and Planning appear less challenging in this context. This overall tendency is also confirmed for our best model. In particular, Figure 3c shows that the model benefits from 32 frames in all categories but Out-of-scope and Uncertainty, where it instead excels with black videos, though in most other cases performance in this setting is poor. Surprisingly, for Counterfactual and Planning, a relatively high accuracy is obtained already with black videos. Interestingly, differences also emerge when comparing the 1-Frame vs. 32-Frames settings. In the majority of cases, the 1-Frame is not enough, while 32 frames increase performance. This difference is less pronounced or does not hold for Planning, Counterfactual and Uncertainty. For example, in the Implicit Total category, the 32-Frames model answers correctly to How does the vehicle move? [^14] Referring to a pirate ship in the amusement park with The vehicle moves in a swinging way, with back-and-forth movements, while the 1-Frame setting hallucinates with The vehicle moves slowly along the amusement park route.
@@
Qwen-2.5-VL 72B 0.62 0.84 0.37 0.65 1.07
Table 4: VLM performance (32-Frames setting) for OEVQA (Task 2) according to similarity-based metrics.
Table 4 highlights how similarity-based metrics (e.g., ROUGE, BLEU) often do not align with semantic correctness. InternVL2 scores highest on surface-level similarity (e.g., BERTScore: $0.84$, CIDEr: $1.18$) – as well as Qwen2.5-VL 72B – but lower when considering the LLM-as-a-judge metric ($0.49$), while Qwen2.5-VL 72B and InternVL3 78B offer more (semantic) accurate answers. LLaVA-NeXT-Video underperforms across all evaluations.
@@ 5.3 Discussion
The results presented for Task 1 and 2 clearly highlight two key findings. First, within each task, the role of the information extracted from videos is unequally distributed across reasoning categories. The star-shaped Figure 3b illustrates such differences: the star’s picks highlight the categories that profit from the visual input the most: Spatial Total, Implicit Total, Causal, Out-of-Scope, and Uncertainty. On the other hand, and quite surprisingly, Qwen2.5-VL 72B handles Out-of-Scope and Uncertainty better when provided the black videos, which are expected to be uninformative, than with the full 32 frames. This calls for a deeper analysis of the reason behind such a result, as the more visual context the model receives, the more it hallucinates, reducing Out-of-Scope scores, and becomes overly assertive, reducing Uncertainty scores. Overall, this shows that MAIA categories are extremely useful to factorize the contribution of the visual vs. linguistic components of VLMs, favoring a more nuanced analysis of their actual abilities. Finally, among the most challenging categories – discussed in 5.1 and 5.2 – several (e.g., Spatial Partial, Implicit Partial, and Temporal Duration) share a temporal dimension, reinforcing the evidence that temporal reasoning still remains a fundamental issue of current models.
A second noteworthy fact is the effect of the task design. By comparing Table 1 and Table 2, we see that overall accuracy is higher in Task $2$ than in Task $1$, even though the underlying information the model has to exploit to perform the tasks is the same, given the alignment between their data points. Such an increase is found across models: +$0.26$ increase on average – in line with the Generative AI Paradox – they are better at generating than at understanding text West et al. (2024). Figure 3b vs. Figure 3c illustrates such a difference for Qwen2.5-VL 72B. Interestingly, the difference is more pronounced for some categories, as shown by the fact that Figure 3c no longer has the shape of a star; for instance, Planning, Counterfactual, Spatial Partial and Temporal Duration improve the most. On the other hand, Uncertainty and Out-of-scope accuracy drops in the NLG task; from a qualitative analysis, we saw that this is mostly due to the generation of hallucinations (Appendix B.3).
However, models’ size proved to be a crucial factor: larger models consistently achieve higher accuracy, even within the same family, suggesting that scaling provides apparent advantages on individual tasks, although this picture changes when considering MAIA’s broader scope (as discussed in the next section).
@@ 6 Aggregating Understanding and Generation
where $M$ is the model, $q$ a question, $S_{q}$ the set of TS-FS pairs, and $a_{M}$ the model’s answers. Intuitively, given a question $q$ on a video $v$, we reward the ability of a VLM to both select a correct answer $TF$ from a TFFSTF-FS pair related to $q$, and to generate a correct answer $a$ to $q$, as discussed in Section 1 when commenting the question What is the cake made of? and the aligned data of the NLU and NLG tasks in Figure 1.
Table 3 reports Agg-Acc for our best model Qwen2.5-VL 72B: $0.47$ with 32-frames, $0.33$ with 1-frame, and only $0.18$ with black video; this shows the more challenging nature of the aggregate task, and that a single frame is not that robust. Planning, Spatial Partial, and Temporal Duration remain challenging even with 32-Frames, despite their notable improvements from Task $1$ to Task $2$. Results confirm that MAIA’s aggregated understanding&generation task creates a harder benchmark, laying the ground toward a more objective VLM evaluation, even when considering larger models that apparently achieve high accuracy and appear to perform well.
@@ 7 Conclusion
We introduce MAIA, a benchmark designed for fine-grained investigation of the reasoning abilities of VLMs on videos. MAIA has two aligned tasks: a visual statement verification task (NLU), and a open-ended visual question answering task (NLG), both on the same set of video related questions. First, we provided a in-depth analysis of the two tasks independently, showing the importance of evaluating model with answer-pools to account for model consistency. Then, we couple comprehension and generation in an aggregated evaluation framework, arguing that the aggregated "all-in-one" understanding&generation task is a challenging and natural setting toward a more objective VLM evaluation, as it reveals inconsistencies within the same task and a lack of robustness across aligned tasks, even in larger models. As for the future, it would be interesting to see whether our framework promote models that undergo learning paradigms tightly integrating these two capabilities, as in Gul and Artzi (2024).
@@
We acknowledge that the number of videos in MAIA may seem relatively small, with only 100 samples. However, their combination with 12 reasoning categories results in 19,200 samples for Task 1 and an equal number of question-answer pairs, providing a robust evaluation set, not intended for any kind of training. Still, these 100 videos constitute only the initial core of a broader evaluation framework planned for future development. Moreover, we are aware that our most probable baseline, constructed using probabilities derived from our set of LLM’s logits, poses a limitation that we intend to address in future work. We plan to compare each VLM with its corresponding LLM to obtain more reliable results for proper comparisons and analyses of potential statistical biases. Regarding Task $2$, we aim to dive deep into the comparison between the similarity metrics used and the emerging topic of LLMs as judges. Additionally, we intend to further investigate this latter approach to assess its actual validity as a reliable evaluation method. This direction will help us refine accuracy metrics, ultimately enhancing our ability to rigorously test model robustness and consistency across our two specular tasks. Furthermore, we acknowledge that our evaluation did not include large-scale proprietary models (e.g., ChatGPT). Our focus was primarily on testing the performance of open-source and easily exploitable language models to provide a comprehensive overview of their capabilities on the benchmark. Finally, we are also aware that our benchmark has not yet been compared with existing ones to assess its relative difficulty and challenge level. This limitation comes from the lack of comparable resources in Italian. As a future direction, we plan to translate MAIA into English and replicate our experiments, enabling more meaningful comparisons with similar English-language benchmarks.